# Second Law of Thermodynamics without Einstein Relation

### v1

@@ Abstract
$\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\int{\text{d}}\mathbf{x}\frac{\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J}}{p}\,\geq\,0\,.$ (19)
Note that the correspondence holds for the means only and not for the full statistics, $\dot{Q}_{t}\neq-k_{\mathrm{B}}\mathcal{T}\dot{\Omega}_{t}$ and $\dot{R}_{t}\neq k_{\mathrm{B}}\mathcal{T}\dot{\Sigma}_{t}$. For example, the variances are different, $\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})p{\text{d}}\mathbf{x}$, whereas $\langle({\text{d}}Q_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int\mathbf{F}\cdot\mathbf{D}\cdot\mathbf{F}p{\text{d}}\mathbf{x}\neq(k_{\mathrm{B}}\mathcal{T})^{2}\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}$. (The subscript $\mathrm{c}$ denotes a cumulant.) Since the Jarzynski equality [35, 36] and the Crooks fluctuation theorem [37] require statistics beyond the mean, they are not restored by the generalized $\mathcal{T}$. Thus, without the ER, the various FRs [34] remain related only to informatic quantities (the IEP and IH) [41] and not to thermodynamics (dissipation and heat).
Another relation to be examined is the one between the heat of stochastic systems and the breaking of FDT. Such a connection was established by Harada and Sasa [59], implicitly assuming the ER and a configuration-independent scalar mobility. (We recall that the ER and FDT are equivalent only in equilibrium [60].) Connections found recently for various active systems between FDT breaking and the IH (e.g., Refs. [39, 40]) have similarly made these two strong assumptions. As discussed above, without these assumptions the IH and the thermodynamic heat are separate quantities. In addition, the FDT is violated [61]. It is then unclear which heat is connected to FDT breaking — the IH or the thermodynamic heat? We find that neither connection can be established in the general case [61], implying that the Harada-Sasa relation breaks down as well.
@@
$\langle\dot{Q}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle\mathbf{F}(\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.-\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle\mathbf{F}(\mathbf{X}_{t},t)\rangle\right)\,.$ (20)
@@
$\langle\dot{\Omega}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.-\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\rangle\right)\,.$ (21)
@@
We now revisit additional classical-thermodynamic results demonstrating the relevance of the generalized temperature $\mathcal{T}$. More details regarding these consequences are found in the SM [61]. (i) Consider a process that starts and ends at the same protocol value, $\Lambda_{t_{\mathrm{f}}}=\Lambda_{0}$, and then allowed to relax to steady state. Since $\oint_{0}^{\infty}{\text{d}}t\langle\dot{S}_{t}\rangle=0$, we obtain the Clausius inequality, $\oint_{0}^{\infty}{\text{d}}t\langle\dot{Q}_{t}\rangle/(k_{\mathrm{B}}\mathcal{T})=-\oint_{0}^{\infty}{\text{d}}t\langle\dot{\Sigma}_{t}\rangle\leq 0$ [61]. (ii) Construct a Carnot-like cyclic engine consisting of an “isothermal” process (with constant $\mathcal{T}_{\mathrm{H}}$ and $\langle Q_{\mathrm{in}}\rangle>0$), an “adiabatic” process (in which the system “cools” down to $\mathcal{T}_{\mathrm{C}}$), followed by another isothermal process (with constant $\mathcal{T}_{\mathrm{C}}$, and $\langle Q_{\mathrm{out}}\rangle<0$), and an adiabatic process (where the system reaches $\mathcal{T}_{\mathrm{H}}$). The efficiency of such an engine is found as $\langle-W_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle=1-\mathcal{T}_{\mathrm{C}}/\mathcal{T}_{\mathrm{H}}+k_{\mathrm{B}}\mathcal{T}_{\mathrm{C}}\langle\Sigma_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle$, where $\langle\Sigma_{\mathrm{cycle}}\rangle$ is the mean IEP during a cycle [61]. Thus, the mean IEP quantifies the cycle’s irreversibility compared to a generalized Carnot efficiency, $1-\mathcal{T}_{\mathrm{C}}/\mathcal{T}_{\mathrm{H}}$. (See also Ref. [62] in the context of nonequilibrium engines.) (iii) Define the nonequilibrium free energy, $A_{t}=H_{t}-k_{\mathrm{B}}\mathcal{T}S_{t}$. It is easy to show that it bounds the extractable work in “isothermal” transformations, $\langle-\dot{W}_{t}\rangle=-(\langle\dot{A}_{t}\rangle+k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle)\leq-\langle\dot{A}_{t}\rangle$ [61].
An interesting consequence of Eq. (15) relates to a particular class of nonequilibrium systems having multiple “temperatures”, e.g., a mixture of different types of particles, each of which is subjected to a different “thermostat” [63, 64]. These models may be relevant, for instance, for mixtures of passive and active particles [65, 66, 67, 68, 69]. They yield a temperature-like scalar which is a weighted average of the thermostats’ temperatures. Considering a diagonal but nonscalar $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$, with different inverse temperatures as its eigenvalues, we may view Eq. (15) as a generalization of such a weighted average. Perturbing different global parameters amounts to taking a perturbation $\mathbf{J}$ along different axes, thus picking up different eigenvalues (inverse temperatures) of $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$. We show a simplistic example in the SM [61].
@@
To restore the information-thermodynamics correspondence we have imposed the second law of thermodynamics directly and identified a generalized temperature $\mathcal{T}$ [Eq. (15)], which gives all the thermodynamically required properties: nonnegative dissipation, zero dissipation at equilibrium, linear thermodynamic stability, and the temperature being a macroscopic (integrated) property. The generalized temperature is model-independent — with a calorimetric measurement of the heat and path-sampling for the IH, one may find the system’s $\mathcal{T}$ with no required knowledge of the underlying dynamics. The variety of tools available for the estimation of IEP from trajectories (e.g., sampling marginal distributions [70, 71, 53] or the thermodynamic uncertainty relations [72, 73]) are particularly useful, since in our construction the average IEP remains connected to the average thermodynamic dissipation. The generalized temperature does not restore the various FRs [35, 37, 34, 59], which in the absence of the ER no longer dictate the statistics of thermodynamic properties [41].
Our choice of generalized temperature in Eq. (15) is not unique. We could have chosen $1/(k_{\mathrm{B}}\mathcal{T})=[\langle\dot{S}_{t}\rangle-f(\mathbf{J})]/\langle\dot{Q}_{t}\rangle$ with any function $f(\mathbf{J})$ satisfying $f(\mathbf{J})\sim J^{2}$ for small $J$ and $f(\mathbf{J}\neq\mathbf{0})>0$. All those temperatures will produce the second law and its consequences and coincide with the conventional temperature $T$ when the ER holds. Our definition of $\mathcal{T}$ stems from the simple quadratic form chosen for the dissipation $R_{t}$, Eq. (19). Beside its simplicity, this choice yields the natural and useful correspondence $\langle\dot{Q}_{t}\rangle=-k_{\mathrm{B}}\mathcal{T}\langle\dot{\Omega}_{t}\rangle$.
The temperature $\mathcal{T}$ as presented here is a property of a single closed system exchanging heat with an arbitrary bath. It dictates the flow of heat into or out of the system based on the change in the system’s entropy. Whether $\mathcal{T}$ functions as an effective temperature in other respects remains to be explored. For example, for two systems in contact (i.e., immersed in a bath that drives each of them out of equilibrium), heat should flow from the system with the higher $\mathcal{T}$ to the one with the lower $\mathcal{T}$. These additional attributes are left for a future study.

### v2

@@ Abstract
$\langle\dot{R}_{t}\rangle=\langle\dot{\Sigma}_{t}\rangle=\int{\text{d}}\mathbf{x}\left(\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J}\right)/p\,\geq\,0\,.$ (19)
Note that the correspondence holds for the means only and not for the full statistics, $\dot{Q}_{t}\neq-k_{\mathrm{B}}\mathcal{T}\dot{\Omega}_{t}$ and $\dot{R}_{t}\neq\dot{\Sigma}_{t}$. For example, the variances of the former are different, $\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})p{\text{d}}\mathbf{x}$, whereas $\langle({\text{d}}Q_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int\mathbf{F}\cdot\mathbf{D}\cdot\mathbf{F}p{\text{d}}\mathbf{x}\neq(k_{\mathrm{B}}\mathcal{T})^{2}\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}$. (The subscript ${\mathrm{c}}$ denotes a cumulant.) Since the Jarzynski equality [37, 38] and the Crooks fluctuation theorem [39] require statistics beyond the mean, they are not restored by the generalized $\mathcal{T}$. Thus, without the ER, the various FRs [36] remain related only to informatic quantities (the IEP and IH) [44] and not to thermodynamics (dissipation and heat).
Another relation to be examined is the one between the heat of stochastic systems and the departure from FDT. Such a connection was established by Harada and Sasa [65], implicitly assuming the ER and a configuration-independent scalar mobility. (We recall that the ER and FDT are equivalent only in equilibrium [66].) Connections found recently for various active systems between FDT violation and the IH (e.g., Refs. [40, 41]) have similarly made these two strong assumptions. As discussed above, without these assumptions the IH and the thermodynamic heat are separate quantities. In addition, the assumptions underlying FDT do not apply [67]. It is then unclear which heat is connected to FDT violation — the IH or the thermodynamic heat? We find that neither connection can be established in the general case [67], implying that the Harada-Sasa relation breaks down as well.
@@
$\langle\dot{Q}_{s}\rangle\!=\!-\!\lim_{t\to s^{+}}\!\left(\frac{{\text{d}}}{{\text{d}}s}\langle\mathbf{F}(\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle+\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle\mathbf{F}(\mathbf{X}_{t},t)\rangle\right)\,.$
@@
$\langle\dot{\Omega}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.+\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\rangle\right)\,.$ (21)
@@
We now revisit additional classical-thermodynamic results demonstrating the relevance of the generalized temperature $\mathcal{T}$. More details regarding these consequences are found in the SM [67]. (i) Consider a process that starts and ends at the same protocol value, $\Lambda_{t_{\mathrm{f}}}=\Lambda_{0}$, and then allowed to relax to steady state. Since $\oint_{0}^{\infty}{\text{d}}t\langle\dot{S}_{t}\rangle=0$, we obtain the Clausius inequality, $\oint_{0}^{\infty}{\text{d}}t\langle\dot{Q}_{t}\rangle/(k_{\mathrm{B}}\mathcal{T})=-\oint_{0}^{\infty}{\text{d}}t\langle\dot{\Sigma}_{t}\rangle\leq 0$ [67]. (ii) Construct a Carnot-like cyclic engine consisting of an “isothermal” process (with constant $\mathcal{T}_{\mathrm{H}}$ and $\langle Q_{\mathrm{in}}\rangle>0$), an “adiabatic” process (in which the system “cools” down to $\mathcal{T}_{\mathrm{C}}$), followed by another isothermal process (with constant $\mathcal{T}_{\mathrm{C}}$, and $\langle Q_{\mathrm{out}}\rangle<0$), and an adiabatic process (where the system reaches $\mathcal{T}_{\mathrm{H}}$). The efficiency of such an engine is found as $\langle-W_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle=1-\mathcal{T}_{\mathrm{C}}/\mathcal{T}_{\mathrm{H}}+k_{\mathrm{B}}\mathcal{T}_{\mathrm{C}}\langle\Sigma_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle$, where $\langle\Sigma_{\mathrm{cycle}}\rangle$ is the mean IEP during a cycle [67]. Thus, the mean IEP quantifies the cycle’s irreversibility compared to a generalized Carnot efficiency, $1-\mathcal{T}_{\mathrm{C}}/\mathcal{T}_{\mathrm{H}}$. (See also Ref. [68] in the context of nonequilibrium engines.) (iii) Define the nonequilibrium free energy, $A_{t}=H_{t}-k_{\mathrm{B}}\mathcal{T}S_{t}$. It is easy to show that it bounds the extractable work in “isothermal” (constant $\mathcal{T}$) transformations, $\langle-\dot{W}_{t}\rangle=-(\langle\dot{A}_{t}\rangle+k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle)\leq-\langle\dot{A}_{t}\rangle$ [67].
An interesting consequence of Eq. (15) relates to a particular class of nonequilibrium systems having multiple “temperatures”, e.g., a mixture of different types of particles, each of which is subjected to a different “thermostat” [69, 70]. These models may be relevant, for instance, for mixtures of passive and active particles [71, 72, 73, 74, 75]. They yield a temperature-like scalar which is a weighted average of the thermostats’ temperatures. Considering a diagonal but nonscalar $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$, with different inverse temperatures as its eigenvalues, we may view Eq. (15) as a generalization of such a weighted average. These systems may attain $\mathbf{J}=\mathbf{0}$, which results in a ‘$/0$ 0’ ambiguity for $\mathcal{T}$. However, perturbing different global parameters amounts to taking a perturbation $\mathbf{J}$ along different axes, thus picking up different eigenvalues (inverse temperatures) of $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$. We show a simplistic example in the SM [67].
@@
To restore the information-thermodynamics correspondence we have imposed the second law of thermodynamics directly and identified a generalized temperature $\mathcal{T}$ [Eq. (15)], which gives all the thermodynamically required properties: nonnegative dissipation, zero dissipation at equilibrium, linear thermodynamic stability, and the temperature being a macroscopic (integrated) property. The generalized temperature is model-independent — with a calorimetric measurement of the heat and path-sampling for the IH, one may find the system’s $\mathcal{T}$ with no required knowledge of the underlying dynamics. In particular, the variety of tools available for the estimation of IEP from trajectories (e.g., sampling marginal distributions [76, 77, 58] or the thermodynamic uncertainty relations [78, 79]) will be useful, since in our construction the average IEP remains connected to the average thermodynamic dissipation. The generalized temperature does not restore the various FRs [37, 39, 36, 65], which in the absence of the ER no longer dictate the statistics of thermodynamic properties [44].
The three conditions which have been used to find $\mathcal{T}$, Eq. (15), do not uniquely determine it. Within linear irreversible thermodynamics [62, 61, 63], they do guarantee a single generalized temperature, as they impose the exact leading $\mathcal{O}(J^{2})$ term, $\langle\dot{R}_{t}\rangle=\int{\text{d}}\mathbf{x}(\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J})/p$. This leading term is the complete dissipation functional when the Einstein relation holds. If one additionally requires this dissipation functional beyond linear theory, $\mathcal{T}$ is then uniquely determined by Eq. (15). The fact that $\mathcal{T}$ is the ratio between $\langle\dot{Q}_{t}\rangle$ (thermodynamics) and $\langle\dot{\Omega}_{t}\rangle$ (information) is then a useful consequence of this choice of dissipation functional.
The temperature $\mathcal{T}$ obtained here is a property of a single closed system exchanging heat with an arbitrary bath. It dictates the direction of heat flow into or out of the system based on the change in the system’s entropy. The fact that the entropy production is non-negative, $\langle\dot{R}_{t}\rangle\geq 0$, implies a positive $\mathcal{T}$ if heat is leaving the system and a negative one if heat is entering the system at the nonequilibrium steady state. At equilibrium, there is no heat flow. It remains to be explored whether $\mathcal{T}$ also dictates the direction of heat flow between two systems in contact.